An investor invented $1700 in two mutual funds. One earned 8% while the other earned 6% profit. If the investor's total profit was $128, how much was invested in each mutual funds.

Answer by checkley79(3341)

.08X+.06(1700-X)=128
.08X+102-.06X=128
.08X-.06X=128-102
.02X=26
X=26/.02
X=1300 AMOUNT INVESTED @ 8%.
PROOF:
.08*1300+.06(1700-1300)=128
104+.06*400=128
104+24=128
128=128